OpenStudy (anonymous):

For the following equations, enter U if the equation has a unique solution, N if it has no solution, and I if it has infinitely many solutions. 4x+9=6 6x+4=6x+9 (1/(4x+9))=(1/(9x+6)) (1/(4x+9))=(2/(4x+9)) (x+4)(x+9)=x^2+13x+36 2(4x+9)=8x+18

OpenStudy (anonymous):

hint :- when x have value then there is unique solution (either real or complex) when x eliminate in the question then there is no solution when u have value like 0=0 or x=x ... ect then there is infinity many solution

OpenStudy (anonymous):

so is it U,N,U,N,U,U? @Halmos

OpenStudy (anonymous):

haha you need to think about the last two equations

OpenStudy (anonymous):

@Halmos isnt it unique because there is a solution? or is it not because x isnt identified

OpenStudy (anonymous):

Never mind I got it, thank you!

OpenStudy (anonymous):

haha wait what did u got ?

OpenStudy (anonymous):

sorry was away a bit

OpenStudy (anonymous):

2(4x+9)=8x+18 8x+18=8x+18 so look at both they are like 0=0 so infinitely many solution same with this (x+4)(x+9)=x^2+13x+36